OpenStudy (anonymous):

At the library, 30 percent of the books checked out are biographies, 50 percent are of the mystery genre, and 20 percent are science fiction. What is the probability that someone checked out wall three kinds of books?

OpenStudy (anonymous):

@AriPotta

OpenStudy (aripotta):

oof. i'm not good with probability, i'm sorry

OpenStudy (anonymous):

It's alright.

OpenStudy (anonymous):

Medal for looking.

OpenStudy (aripotta):

@Hero

OpenStudy (anonymous):

Please help me.

OpenStudy (anonymous):

sorry i haven't done probability in years

OpenStudy (anonymous):

It's okay...

OpenStudy (anonymous):

@Kitt020912

OpenStudy (anonymous):

I got you bro, I got you(:

OpenStudy (anonymous):

thanks;)

OpenStudy (anonymous):

wow ur a math genius.

OpenStudy (anonymous):

are you missing any info in this problem? like the number of books?

OpenStudy (anonymous):

no... thats what it says.

OpenStudy (anonymous):

look, I'll attach it straight from the sheet.

OpenStudy (anonymous):

one sec.

OpenStudy (anonymous):

k

OpenStudy (anonymous):

Mhm:)

OpenStudy (anonymous):

do you know if the want a fraction or..?

OpenStudy (anonymous):

Probably the fraction,

pooja195 (pooja195):

im not sure

OpenStudy (anonymous):

\[\frac{ 30 }{ 100 }\rightarrow biography\] \[\frac{ 50 }{ 100 }\rightarrow mystery\] \[\frac{ 20 }{ 100 }\rightarrow science\] multiply all that together.. \[\frac{ 30 }{ 100 }\times \frac{ 50 }{ 100 }\times \frac{ 20 }{ 100 }=\frac{ 30 }{ 100 }\] and if you had to change\[\frac{ 30 }{ 100 }\]it would be \[0.3\] ☺

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

let me read it all,k?

OpenStudy (anonymous):

OMG I get it